# TabArena: A Living Benchmark for Machine Learning on Tabular Data

### v2

@@ Results.
Figure A.1 presents a leaderboard including all models. We impute the results for models on datasets where they are not applicable with the results of RandomForest (default). We choose the default random forest since it is a fast baseline that is sufficiently but not unreasonably weak, to penalize models that are not applicable to all datasets. Table A.1 presents the same data in tabularized format, akin to the current version of the live leaderboard at tabarena.ai. Table A.1 further includes several additional metrics to asses peak average performance, some of which change the ranking (see the color coding) as they are less influenced by model-wise negative outlier results introduced by imputation.
We further investigate our results by presenting the leaderboard across task types. We show the results per task type by computing the results only with datasets from: binary classification in Figure A.2, multiclass classification in Figure A.3, and regression in Figure A.4.
Next, Figure A.5 and Figure A.6 present the results for the TabPFNv2-compatible and TabICL-compatible datasets, but also impute TabPFNv2/TabICL to enable a more direct comparison between these two foundation models. Finally, Figure A.7 presents the result only with datasets for which both TabPFNv2 and TabICL are compatible.
@@
RealMLP (T+E) 1569-27,+26 0.537 8.5 4.9 1 6.4% 6564.71 10.26
TabM (T+E) 1552-29,+21 0.502 9.2 4.4 4 6.7% 3285.87 1.47
LightGBM (T+E) 1532-26,+29 0.452 9.9 5.5 2 8.3% 416.98 2.64
CatBoost (T+E) 1489-23,+25 0.433 11.6 7.3 0 7.6% 1658.41 0.65
CatBoost (T) 1475-27,+22 0.416 12.2 6.1 2 7.8% 1658.41 0.08
TabM (T) 1455-34,+22 0.413 13.0 6.7 1 7.7% 3285.87 0.17
LightGBM (T) 1450-26,+27 0.363 13.1 10.7 0 9.0% 416.98 0.33
XGBoost (T+E) 1446-26,+23 0.361 13.4 8.1 1 9.1% 693.49 1.69
ModernNCA (T+E) 1435-28,+24 0.444 13.9 5.0 3 8.4% 4621.67 8.15
CatBoost (D) 1432-24,+27 0.372 14.0 7.0 2 9.0% 6.83 0.08
TabPFNv2 (T+E) 1414-26,+25 0.476 14.9 3.1 11 8.3% 3030.15 21.44
XGBoost (T) 1409-20,+24 0.313 15.0 12.1 0 9.4% 693.49 0.31
ModernNCA (T) 1404-30,+31 0.307 15.3 8.6 1 9.1% 4621.67 0.47
TabICL (D) 1388-24,+27 0.381 16.0 4.5 6 8.9% 6.63 1.48
RealMLP (T) 1349-21,+26 0.229 17.8 15.1 0 9.8% 6564.71 0.49
TabM (D) 1348-20,+28 0.285 17.8 11.9 0 10.7% 10.49 0.13
TabPFNv2 (T) 1346-23,+22 0.360 18.0 5.6 1 10.4% 3030.15 0.46
TorchMLP (T+E) 1334-18,+29 0.220 18.5 13.4 0 10.3% 2874.67 1.95
TabPFNv2 (D) 1317-27,+28 0.324 19.4 5.5 4 11.4% 3.36 0.31
ModernNCA (D) 1302-23,+23 0.159 20.2 12.2 1 12.6% 14.87 0.31
TabDPT (D) 1298-21,+28 0.300 20.3 4.8 7 12.0% 22.53 8.55
EBM (T+E) 1292-23,+25 0.184 20.6 11.6 1 14.4% 1331.68 0.20
FastaiMLP (T+E) 1250-21,+25 0.158 22.6 13.3 0 13.7% 593.24 4.47
RealMLP (D) 1246-22,+29 0.104 22.9 19.4 0 12.4% 21.86 0.84
ExtraTrees (T+E) 1242-28,+27 0.124 23.0 15.0 0 14.1% 183.02 0.76
EBM (T) 1236-25,+25 0.135 23.3 16.2 0 15.0% 1331.68 0.02
XGBoost (D) 1229-23,+29 0.115 23.5 18.3 0 12.4% 1.94 0.12
TorchMLP (T) 1223-28,+25 0.113 23.8 20.2 0 12.4% 2874.67 0.13
LightGBM (D) 1206-28,+28 0.088 24.8 22.0 0 13.3% 1.96 0.14
EBM (D) 1202-30,+24 0.134 24.9 13.2 1 16.0% 4.67 0.04
RandomForest (T+E) 1201-23,+23 0.098 25.0 12.9 1 14.9% 373.18 0.77
ExtraTrees (T) 1198-28,+22 0.094 25.1 16.6 0 15.1% 183.02 0.09
FastaiMLP (T) 1159-20,+22 0.073 26.8 21.5 0 15.4% 593.24 0.31
RandomForest (T) 1150-26,+25 0.072 27.2 16.6 0 15.9% 373.18 0.09
TorchMLP (D) 1067-26,+24 0.022 30.7 28.2 0 17.2% 9.99 0.13
FastaiMLP (D) 1009-24,+31 0.023 32.7 29.9 0 20.6% 2.86 0.37
RandomForest (D) 1000-0,+0 0.009 33.1 31.3 0 21.0% 0.43 0.05
ExtraTrees (D) 970-23,+36 0.025 34.0 30.3 0 22.8% 0.25 0.05
Linear (T+E) 919-28,+29 0.031 35.6 27.6 0 30.6% 47.49 0.17
Linear (T) 882-28,+23 0.020 36.6 31.9 0 31.4% 47.49 0.07
Linear (D) 856-29,+36 0.014 37.1 29.9 0 32.7% 1.52 0.09
KNN (T+E) 684-27,+23 0.000 40.5 40.2 0 45.3% 2.74 0.18
KNN (T) 607-40,+33 0.000 41.4 41.3 0 46.8% 2.74 0.04
KNN (D) 456-46,+47 0.000 42.8 42.6 0 54.1% 0.05 0.02
@@ A.2 Analyzing Training Time Limit
In our experiments, we restrict the time to evaluate one configuration on one train split of a dataset to $1$ hour. Thus, a model must finish training (across all $8$ inner folds) within $1$ hour, or its training will be gracefully stopped early. Figure A.8 presents the training runtime for all hyperparameter configurations for all models by visualizing what proportion of configurations (x-axis) took how many seconds for training (y-axis).
We observe that for all models, less than $1\%$ of all configurations reach the time limit of $1$ hour. We further investigate the time limit for the GPU-optimized models in Section A.3. For EBMs, we notice that the training was not stopped early at the $1$ hour time limit, positively influencing its results. As this only concerns a small fraction of hyperparameter trials, we did not rerun the training for EBM.
@@
In our main experiments, we ran TabM and ModernNCA on GPU. To further investigate the impact of hardware choice on these models, we also ran TabM and ModernNCA on CPU. Moreover, we ran RealMLP on GPU. Figure A.9 demonstrates that the hyperparameter configurations of TabM and ModernNCA train much faster and RealMLP slightly faster on GPU than on CPU. We conclude from this ablation that training TabM and ModernNCA on CPU with a time limit of $1$ hour would negatively influence their predictive performance. While the influence is marginal for RealMLP, it seems non-marginal for TabM and ModernNCA.
Figure A.9: Runtime Analysis of RealMLP, TabM, and ModernNCA on CPU vs. GPU. We show the training runtime distribution of RealMLP, TabM, and ModernNCA trained on CPU and GPU across all TabArena datasets. For ModernNCA and TabM on CPU, approximately $16\%$ of runs are early stopped due to the $1$-hour time limit. For GPU, less than $0.1\%$ of runs are early stopped due to the time limit.
@@ A.4 TabArena-Lite
Benchmarking can quickly become very expensive, especially with a sophisticated protocol to guarantee robust results. To reduce the cost of benchmarking, we introduce TabArena-Lite. TabArena-Lite is a continually maintained subset of TabArena that consists, in its first version, of all datasets with one outer fold. Figure A.10 shows results on TabArena-Lite, using 200 hyperparameter configurations per model, but only a single outer fold for all datasets. The results are similar to the results on TabArena in Figure 1, showing that TabArena-Lite is a good indicator of model performance.
@@ A.5 Investigating Statistical Significance
We investigate the statistical significance between models by using critical difference diagrams (CDDs) to represent the results of a Friedman test and then a Nemenyi post-hoc test ($\alpha=0.05$) from AutoRank [^1] https://github.com/sherbold/autorank. Figures A.11, A.12 and A.13 show the CDDs for the full benchmark, TabPFNv2-compatible datasets, and TabICL-compatible datasets with respect to the peak performance of the models, i.e., tuned + ensembled where available. We further investigate statistical significance per-dataset in Appendix E.
@@ Appendix B Data curation
For initializing the TabArena benchmark, we surveyed all the datasets used in 13 previous benchmarking studies: $450$ from PMLB(Mini) [43, 44, 45], $72$ from OpenML-CC18 [29], $45$ from Grinsztajn et al. [32], $11$ from Shwartz-Ziv and Armon [31], $11$ from Gorishniy et al. [30], $176$ from TabZilla [33], $35$ from OpenML-CTR23 [34], $104$ from AMLB [35], $8$ from TabRed [10], $10$ from Tschalzev [8], $279$ from TabRepo [37], $118$ from PyTabKit [20], and $300$ from TALENT [47, 36].
These studies were selected with the goal of covering a wide range of datasets used in tabular benchmarking so that we can clean up the field from problematic or unsuitable datasets. Therefore, each of the studies represents a frequently used benchmark or a general milestone study in the field of tabular machine learning. Combining the dataset collections results in 1053 uniquely named datasets.
For TabArena-v0.1, we aimed at using only those datasets representing realistic, predictive tabular data tasks practitioners would be interested to solve. Therefore, we define a set of selection criteria described in Section B.1. Two of the coauthors manually investigated each of the datasets and applied our selection criteria. We publicly share their notes and curated metadata: tabarena.ai/dataset-curation. Furthermore, we share insights from our curation process in Section B.3
Importantly, we did not exhaustively test each dataset for each of our curation criteria, but proceeded with the next dataset whenever a dataset clearly met at least one of our criteria for exclusion. Therefore, Figure 2 represents the first reason for exclusion that we noticed in a dataset.
Surprisingly, only 51 datasets satisfying all criteria remained. Section B.2 provides additional information on the selected datasets. We consider our data curation a clean-up for tabular data benchmarking that is necessary, but imperfect. Therefore, we aim to continuously improve the data selection and invite researchers to challenge our documented decisions. Section D.3 details protocols to contribute new datasets to TabArena by applying our criteria.
@@ B.1 Dataset Selection Criteria
We exclude datasets that suffer from one of the following data quality issues: (1) heavily preprocessed datasets, such as those where the whole dataset was already used for preprocessing in a way that leaks the target (i.e., PCA); (2) datasets for which we could not find sufficient information to judge their source and preprocessing; (3) datasets with an irreversible target leak. In general, we try to find the original state of the dataset and include it, if applicable. We do not generally exclude preprocessed datasets, as datasets are rarely published without any preprocessing, e.g., due to anonymization. We leave a benchmark with model-specific, domain-specific pre-processing per dataset for future work.
@@
We exclude datasets that cannot be made to be automatically downloaded from a tabular data repository. Eligible data repositories must be open-access, i.e., users do not need an account to download data. Furthermore, the repositories require a structured data and task representation, including metadata information such as feature types, the target column, and outer splits. To the best of our knowledge, only OpenML fulfills these requirements so far. If applicable due to licensing, we manually upload datasets to OpenML to include them in TabArena. This criterion is necessary to enable automated benchmarking and a straightforward user experience.

### v3

@@ Results.
Figure˜A.1 presents a leaderboard including all models. We impute the results for models on datasets where they are not applicable with the results of RandomForest (default). We choose the default random forest since it is a fast baseline that is sufficiently but not unreasonably weak, to penalize models that are not applicable to all datasets. Table˜A.1 presents the same data in tabularized format, akin to the current version of the live leaderboard at tabarena.ai. Table˜A.1 further includes several additional metrics to asses peak average performance, some of which change the ranking (see the color coding) as they are less influenced by model-wise negative outlier results introduced by imputation.
We further investigate our results by presenting the leaderboard across task types. We show the results per task type by computing the results only with datasets from: binary classification in Figure˜A.2, multiclass classification in Figure˜A.3, and regression in Figure˜A.4.
In addition, we show pairwise win rate comparisons of the models in Figure˜A.5. Among the models with the best performance in our main analysis, the win rates over strong competitors are rather moderate. This undermines that there is no one-size-fits-all solution.
Next, Figure˜A.6 and Figure˜A.7 present the results for the TabPFNv2-compatible and TabICL-compatible datasets, but also impute TabPFNv2/TabICL to enable a more direct comparison between these two foundation models. Finally, Figure˜A.8 presents the result only with datasets for which both TabPFNv2 and TabICL are compatible.
@@
RealMLP (T+E) 1564-57,+73 0.532 8.6 4.9 1.0 6.4% 6564.71 10.26
TabM (T+E) 1541-61,+74 0.499 9.4 4.4 4.0 6.7% 3285.87 1.47
LightGBM (T+E) 1527-45,+49 0.448 9.9 5.5 2.0 8.3% 416.98 2.64
CatBoost (T+E) 1482-55,+57 0.428 11.7 7.3 0.0 7.6% 1658.41 0.65
CatBoost (T) 1469-49,+56 0.411 12.2 6.1 2.0 7.8% 1658.41 0.08
TabM (T) 1447-52,+73 0.410 13.2 6.7 1.0 7.7% 3285.87 0.17
LightGBM (T) 1446-37,+39 0.359 13.2 10.7 0.0 9.0% 416.98 0.33
XGBoost (T+E) 1440-47,+50 0.358 13.5 8.2 1.0 9.1% 693.49 1.69
CatBoost (D) 1427-43,+52 0.369 14.1 7.0 2.0 9.0% 6.83 0.08
ModernNCA (T+E) 1425-67,+103 0.444 14.2 5.0 3.0 8.4% 4621.67 8.15
TabPFNv2 (T+E) 1405-76,+83 0.474 15.1 3.1 11.0 8.3% 3030.15 21.44
XGBoost (T) 1404-44,+49 0.309 15.1 12.2 0.0 9.4% 693.49 0.31
ModernNCA (T) 1396-44,+55 0.305 15.5 8.6 1.0 9.1% 4621.67 0.47
TabICL (D) 1383-64,+71 0.380 16.1 4.5 6.0 8.9% 6.63 1.48
TabM (D) 1343-55,+68 0.282 18.0 11.9 0.0 10.7% 10.49 0.13
RealMLP (T) 1342-45,+43 0.226 18.1 15.3 0.0 9.8% 6564.71 0.49
TabPFNv2 (T) 1336-72,+77 0.359 18.4 5.7 1.0 10.4% 3030.15 0.46
TorchMLP (T+E) 1327-53,+49 0.215 18.8 13.6 0.0 10.3% 2874.67 1.95
TabPFNv2 (D) 1309-81,+87 0.323 19.7 5.5 4.0 11.4% 3.36 0.31
ModernNCA (D) 1294-48,+58 0.158 20.5 12.3 1.0 12.6% 14.87 0.31
TabDPT (D) 1290-72,+79 0.298 20.7 4.8 7.0 12.0% 22.53 8.55
EBM (T+E) 1286-63,+59 0.182 20.8 11.6 1.0 14.4% 1331.68 0.20
FastaiMLP (T+E) 1242-64,+64 0.156 23.0 13.4 0.0 13.7% 593.24 4.47
RealMLP (D) 1238-47,+51 0.103 23.2 19.6 0.0 12.4% 21.86 0.84
ExtraTrees (T+E) 1236-50,+54 0.123 23.4 15.1 0.0 14.1% 183.02 0.76
EBM (T) 1231-55,+55 0.132 23.6 16.3 0.0 15.0% 1331.68 0.02
XGBoost (D) 1225-50,+49 0.115 23.9 18.3 0.0 12.4% 1.94 0.12
TorchMLP (T) 1220-54,+59 0.111 24.1 20.3 0.0 12.4% 2874.67 0.13
EBM (D) 1198-66,+64 0.133 25.2 13.2 1.0 16.0% 4.67 0.04
RandomForest (T+E) 1197-56,+62 0.096 25.2 13.0 1.0 14.9% 373.18 0.77
LightGBM (D) 1197-45,+44 0.088 25.3 22.3 0.0 13.3% 1.96 0.14
ExtraTrees (T) 1190-52,+55 0.093 25.6 17.7 0.0 15.1% 183.02 0.09
FastaiMLP (T) 1154-61,+60 0.070 27.3 21.8 0.0 15.4% 593.24 0.31
RandomForest (T) 1149-49,+59 0.071 27.5 15.1 1.0 15.9% 373.18 0.09
TorchMLP (D) 1065-49,+50 0.022 31.2 28.5 0.0 17.2% 9.99 0.13
FastaiMLP (D) 1007-71,+66 0.022 33.4 30.4 0.0 20.6% 2.86 0.37
RandomForest (D) 1000-44,+50 0.009 33.7 31.8 0.0 21.0% 0.43 0.05
KNN (T+E) 984-79,+66 0.027 34.2 30.7 0.0 23.3% 129.01 1.80
ExtraTrees (D) 980-75,+72 0.024 34.3 30.6 0.0 22.8% 0.25 0.05
Linear (T+E) 967-110,+72 0.036 34.8 26.8 0.0 28.9% 237.58 0.42
Linear (T) 936-115,+72 0.027 35.8 30.1 0.0 29.7% 237.58 0.08
KNN (T) 858-93,+64 0.011 38.0 36.3 0.0 28.8% 129.01 0.18
Linear (D) 836-136,+84 0.014 38.5 30.8 0.0 32.7% 1.19 0.12
KNN (D) 614-133,+102 0.000 42.2 41.7 0.0 43.0% 0.19 0.04
@@
Figure A.5: Pairwise win rate comparison for all datasets. Higher numbers correspond to a better win rate for the model on the y-axis.
@@ A.2 Analyzing Training Time Limit
In our experiments, we restrict the time to evaluate one configuration on one train split of a dataset to $1$ hour. Thus, a model must finish training (across all $8$ inner folds) within $1$ hour, or its training will be gracefully stopped early. Figure˜A.9 presents the training runtime for all hyperparameter configurations for all models by visualizing what proportion of configurations (x-axis) took how many seconds for training (y-axis).
We observe that for all models, less than $1\%$ of all configurations reach the time limit of $1$ hour. We further investigate the time limit for the GPU-optimized models in Section˜A.3. For EBMs, we notice that the training was not stopped early at the $1$ hour time limit, positively influencing its results. As this only concerns a small fraction of hyperparameter trials, we did not rerun the training for EBM.
@@
In our main experiments, we ran TabM and ModernNCA on GPU. To further investigate the impact of hardware choice on these models, we also ran TabM and ModernNCA on CPU. Moreover, we ran RealMLP on GPU. Figure˜A.10 demonstrates that the hyperparameter configurations of TabM and ModernNCA train much faster and RealMLP slightly faster on GPU than on CPU. We conclude from this ablation that training TabM and ModernNCA on CPU with a time limit of $1$ hour would negatively influence their predictive performance. While the influence is marginal for RealMLP, it seems non-marginal for TabM and ModernNCA.
Figure A.10: Runtime Analysis of RealMLP, TabM, and ModernNCA on CPU vs. GPU. We show the training runtime distribution of RealMLP, TabM, and ModernNCA trained on CPU and GPU across all TabArena datasets. For ModernNCA and TabM on CPU, approximately $16\%$ of runs are early stopped due to the 11-hour time limit. For GPU, less than $0.1\%$ of runs are early stopped due to the time limit.
@@ A.4 TabArena-Lite
Benchmarking can quickly become very expensive, especially with a sophisticated protocol to guarantee robust results. To reduce the cost of benchmarking, we introduce TabArena-Lite. TabArena-Lite is a continually maintained subset of TabArena that consists, in its first version, of all datasets with one outer fold. Figure˜A.11 shows results on TabArena-Lite, using 200 hyperparameter configurations per model, but only a single outer fold for all datasets. The results are similar to the results on TabArena in Figure˜1, showing that TabArena-Lite is a good indicator of model performance.
@@ A.5 Investigating Statistical Significance
We investigate the statistical significance between models by using critical difference diagrams (CDDs) [58] to represent the results of a Friedman test and then a Nemenyi post-hoc test ($\alpha=0.05$) from AutoRank [^1] https://github.com/sherbold/autorank [59]. Figures˜A.12, A.13 and A.14 show the CDDs for the full benchmark, TabPFNv2-compatible datasets, and TabICL-compatible datasets with respect to the peak performance of the models, i.e., tuned + ensembled where available. We further investigate statistical significance per-dataset in Appendix˜F.
@@
A.6 Investigating Tuning Trajectories and Validation Overfitting
In Figure˜5 (right), most models appear to have saturated their performance by 201 tuning configs, indicating that further increasing the amount of configurations beyond 201 is unlikely to lead to significant improvements without also expanding the model search spaces. ModernNCA’s performance peaks at 25 configurations and then progressively degrades with further tuning, which can be attributed to validation data overfitting during tuning; also often called overtuning [41, 42]. This can be observed in Figure˜A.15, where despite ModernNCA being the 6th place model in terms of Elo when evaluated on the test data, it is the 1st place model when evaluated on the validation data. In other words, we observe that overtuning occurs since overfitting on the validation data results in reduced test performance, despite continued improvement in the validation score. Furthermore, the degree of overfitting for each method varies widely. As seen in Figure˜A.16, neural network models overfit much more than other methods, even with far fewer configurations. ModernNCA, in particular, only needs 10 random configurations ensembled together to overfit more than any non-deep learning method with 201 ensembled configurations. TabM stands out as the sole exception, demonstrating remarkable resilience to overfitting, possibly due to its internal ensembling, joint early stopping for the whole ensemble, and moderately-sized hyperparameter search space.
With these insights, we can now explain why the TabArena ensemble overwhelmingly favors ModernNCA and RealMLP over other models (Figure˜7). The reason is not necessarily because these models are the correct ones to choose, but rather because the ensemble is optimizing the validation performance as a proxy for the test performance, and the models that are most overfit on the validation data will naturally be over-selected by the ensembling algorithm (and likewise for model selection). Similarly, models that perform very well on test data while avoiding overfitting (such as TabM) are heavily under-selected by the ensemble. Despite these issues, the TabArena ensemble still drastically outperforms all methods, including AutoGluon. This indicates that a novel overfitting-aware method for selecting configurations in the TabArena ensemble could yield significantly improved performance, further advancing the state of the art. We leave the exploration of such algorithms and other solutions for overtuning to future work.
Figure A.15: The tuning trajectories of model ensembles for (Left) Elo when evaluated on the test data. (Right) Elo when evaluated on the validation data. Time is shown as the tuning time with points from left to right marking ensembles of increasing numbers of random configurations (1, 2, 5, 10, 25, 50, 100, 150, 201). The trajectories are sampled 20 times from all trials and averaged. The right-most highlighted points use all configurations. The left figure calibrates 1000 Elo to the test performance of the default random forest configuration. The right figure calibrates 1000 Elo to the validation performance of the default random forest configuration.
Figure A.16: The validation overfitting tuning trajectories of model ensembles for Elo when evaluated on the test data vs. Elo when evaluated on the validation data. The setup is identical to Figure˜A.15. A value of 800 on the y-axis means that the configuration has 800 higher elo based on its validation scores than on its test scores relative to default random forest. Higher values lead to less reliable estimates of a model’s performance on test data given only its validation scores, and make effective model selection and post-hoc ensembling more challenging.
@@ Appendix B Data Curation
For initializing the TabArena benchmark, we surveyed all the datasets used in 14 previous benchmarking studies: $450$ from PMLB(Mini) [46, 47, 48], $72$ from OpenML-CC18 [29], $45$ from Grinsztajn et al. [32], $11$ from Gorishniy et al. [30], $11$ from Shwartz-Ziv and Armon [31], $176$ from TabZilla [33], $35$ from OpenML-CTR23 [34], $104$ from AMLB [35], $187$ from Kohli et al. [7], 88 from TabRed [10], $10$ from Tschalzev et al. [8], $279$ from TabRepo [37], $118$ from PyTabKit [20], and $300$ from TALENT [50, 36, 60].
These studies were selected with the goal of covering a wide range of datasets used in tabular benchmarking so that we can clean up the field from problematic or unsuitable datasets. Therefore, each of the studies represents a frequently used benchmark or a general milestone study in the field of tabular machine learning. Note that PMLB(Mini) [46, 47, 48] is not included in table 2 as the reference publication did not include an evaluation of methods. Combining the dataset collections results in 1053 uniquely named datasets.
For TabArena-v0.1, we aimed at using only those datasets representing realistic, predictive tabular data tasks that practitioners would be interested in solving. Therefore, we define a set of selection criteria described in Section˜B.1. Two of the coauthors manually investigated each of the datasets and applied our selection criteria. We publicly share their notes and curated metadata: tabarena.ai/dataset-curation. Furthermore, we share insights from our curation process in Section˜B.3
Importantly, we did not exhaustively test each dataset for each of our curation criteria, but proceeded with the next dataset whenever a dataset clearly met at least one of our criteria for exclusion. Therefore, Figure˜2 represents the first reason for exclusion that we noticed in a dataset.
Surprisingly, only 51 datasets satisfying all criteria remained. Section˜B.2 provides additional information on the selected datasets. Moreover, we share all tasks and datasets as an OpenML suite [29, 61, 62] (ID $457$, alias "tabarena-v0.1"). We consider our data curation a clean-up for tabular data benchmarking that is necessary, but imperfect. Therefore, we aim to continuously improve the data selection and invite researchers to challenge our documented decisions. Section˜E.3 details protocols to contribute new datasets to TabArena by applying our criteria.
@@ B.1 Dataset Selection Criteria
We exclude datasets that suffer from one of the following data quality issues: (1) heavily preprocessed datasets, such as those where the whole dataset was already used for preprocessing in a way that leaks the target or any information about the feature distribution of the test set (e.g., PCA features computed over all data points); (2) datasets for which we could not find sufficient information to judge their source and preprocessing; (3) datasets with an irreversible target leak. In general, we try to find the original state of the dataset and include it, if applicable. We do not generally exclude preprocessed datasets, as datasets are rarely published without any preprocessing, e.g., due to anonymization. We leave a benchmark with model-specific, domain-specific pre-processing per dataset for future work.
@@
We exclude datasets that cannot be made to be automatically downloaded from a tabular data repository. Eligible data repositories must be open-access, i.e., users do not need an account to download data. Furthermore, the repositories require a structured data and task representation, including metadata information such as feature types, the target column, and outer splits. To the best of our knowledge, only OpenML [62] fulfills these requirements so far. If applicable due to licensing, we manually upload datasets to OpenML to include them in TabArena. This criterion is necessary to enable automated benchmarking and a straightforward user experience.